# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ IV.2 Constant of Wente inequality for a metric and consequence
Since then, the study of this structure has been intensively studied. Bethuel–Ghidaglia [20] proved that the constant $C$ is independent of the domain. Ge [63], Braraket [11] and Topping [184] studied such equations on surfaces and proved that the constant are independent of the underlying surface. A weighted Wente inequality has also been studied by Da Lio, Gianocca and Rivière in [46, 65]. In this section, we provide a new proof of the independence of the constant with respect to the underlying metric and domain in Proposition IV.8. As an application, we obtain the existence of a Coulomb frame under the smallness of the Gauss curvature in Proposition IV.10.
@@ IV.2.1 Wente inequality
Let $\Omega\subset\mathbb{R}^{2}$ be a smooth simply connected open set. Let $g$ be a smooth metric on $\Omega$. Let $a,b\in C^{\infty}(\Omega)$ and $u\in C^{\infty}(\Omega)$ be the solution to
$\left\{\begin{aligned} &-\Delta_{g}u=*_{g}(da\wedge db)\quad&&\text{in }\Omega,\\ &u=0\quad&&\text{on }\partial\Omega.\end{aligned}\right$. (IV.15)
@@
  • The first equation of (IV.15) can also be written as
@@
By elliptic regularity, it holds $G_{p}\in C^{\infty}(\Omega\setminus\{p\})$. Thanks to Sard’s theorem (see for instance [108, Chapter 6, Theorem 6.10]), for almost every $\alpha>0$, the set $G_{p}^{-1}(\{\alpha\})$ is a regular curve in $\Omega$. We orient this curve with the vector field
@@
$\int_{\omega_{p}(\alpha,\beta)}|dG_{p}|^{2}_{g}\,d\textup{vol}_{g}=\int_{\alpha}^{\beta}\int_{G_{p}^{-1}(\{s\})}|dG_{p}|_{g}\,d\textup{vol}_{g}\,ds=\beta-\alpha.$ (IV.16)
Now, the solution $u$ to (IV.15) is given by
$u(p)=\int_{\Omega}G_{p}(x)\,da\wedge db.$ (IV.17)
@@
$\leq 2\int_{\omega_{p}(n,n+1)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}.$ (IV.18)
From (IV.17), it holds
@@
$=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$ (IV.19)
$\qquad+\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}.$ (IV.20)
We estimate the term (IV.19) by Cauchy–Schwarz:
@@
$\leq 2\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}.$ (IV.21)
We estimate the term (IV.20) by integration by parts:
@@
$=$ $-\sum_{n\in\mathbb{N}^{*}}\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}.$ (IV.22)
@@
Coming back to (IV.22), we obtain
@@
Using the choice of the $\alpha_{n}$ in (IV.18), we obtain
@@
$\leq 8\,\|da\|_{L^{2}(\Omega,g)}^{2}+8\|db\|_{L^{2}(\Omega,g)}^{2}=16.$ (IV.23)
We used Cauchy–Schwarz and the equalities $\|da\|_{L^{2}(\Omega,g)}=\|db\|_{L^{2}(\Omega,g)}=1$ for the last estimate. Coming back to (IV.17), together with (IV.21) and (IV.23), we deduce that
@@
We multiply (IV.15) by $u$ and integrate by parts:
$\int_{\Omega}|du|^{2}_{g}\,d\textup{vol}_{g}$ $=-\int_{\Omega}u\,\Delta_{g}u\,d\textup{vol}_{g}$
$\leq\|u\|_{L^{\infty}(\Omega)}\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}\leq 18.$
@@ IV.2.2 Estimate of the moving frame
Thanks to [80, Lemma 4.1.3], there exists a frame $(\vec{e}_{1},\vec{e}_{2})\in W^{1,2}(\Omega,g)$ such that
$\begin{cases}d^{*_{g}}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{in }\Omega,\\ \left\langle\partial_{\nu}\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{on }\partial\Omega,\end{cases}$ (IV.24)
where $\nu$ is the unit outward-pointing normal for $g$. Moreover, by decomposition each $d\vec{e}_{i}$ in the sum of its tangential part along $\vec{\Phi}$ and its normal part, we obtain the following estimate
$\int_{\Omega}|d\vec{e}_{1}|^{2}_{g,\mathbb{R}^{m}}+|d\vec{e}_{2}|^{2}_{g,\mathbb{R}^{m}}\,d\textup{vol}_{g}\leq 2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$ (IV.25)
By Poincaré Lemma and (IV.24), there exists $\lambda\in W^{1,2}(\Omega,g)$ such that
$\begin{cases}d\lambda=*_{g}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.26)
Indeed, the boundary condition in (IV.24) implies that $\lambda$ is constant on $\partial\Omega$. Since the first relation of (IV.26) depends only on $d\lambda$ and not on $\lambda$ itself, we can add a constant in order to get the second relation of (IV.26). We apply $d^{*_{g}}$ to (IV.26) to obtain
$\left\{\begin{aligned} &-\Delta_{g}\lambda=-*_{g}d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)\quad\text{in }\Omega,\\ &\lambda=0\quad\text{on }\partial\Omega.\end{aligned}\right$. (IV.27)
@@
Denote $\varepsilon_{0}\coloneqq\|K_{g}\|_{L^{1}(\Omega,g)}$. We multiply the first relation of (IV.27) by $\lambda$ and integrate by parts:
@@ IV.3 Local injectivity and integral densities
In particular, $\nabla f\in VMO(B^{n})$. Then there exist a constant $C>0$ and an open neighborhood $U$ of the origin, both depending only on $\Lambda,\,\omega,\,n,\,m$ such that
@@
$|Mv|\geq C_{0}|v|.$ (IV.28)
Moreover, $C_{0},\,\varepsilon$, and $K$ can be chosen to depend on $\Lambda,\,\omega,\,n,\,m$ only. We choose $r_{0},\varepsilon_{0}\in(0,1/8)$ small enough (to be determined later) depending on $\Lambda,\,\omega,\,n,\,m$ only such that for every ball $B$ of radius $\leq 4r_{0}$ we have $\fint_{B}\left|\nabla f-\fint_{B}\nabla f\right|<\min(\varepsilon,\varepsilon_{0})$. Now fix $x_{0},y_{0}\in B_{r_{0}}(0)$, we denote $r\coloneqq|x_{0}-y_{0}|$, $B^{\prime}\coloneqq B_{2r}(x_{0})$, and
@@
$\fint_{B^{\prime}}|\nabla f-M_{0}|<\min(\varepsilon,\varepsilon_{0}),$
and hence $M_{0}\in K^{\prime}$. Now let $f_{0}(x)\coloneqq f(x)-M_{0}x$, then apply Morrey’s inequality [58, Theorem 5.6.4] in the case $p=2n$ and John–Nirenberg inequality [77, Theorem 3.5], for $\varepsilon_{0}$ small enough we have
$\begin{aligned} |f_{0}(x_{0})-f_{0}(y_{0})|&\leq C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f_{0}|^{2n}\right)^{\frac{1}{2n}}\\ &=C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f-M_{0}|^{2n}\right)^{\frac{1}{2n}}\\[5.69054pt] &\leq C_{2}\mkern 2.0mur\mkern 2.0mu\varepsilon_{0}=2^{-1}C_{0}\mkern 2.0mur\\[8.53581pt] &=2^{-1}C_{0}|x_{0}-y_{0}|,\end{aligned}$ (IV.29)
where $C_{1},\,C_{2}$ are positive constants depending only on $n$, and we choose $\varepsilon_{0}=(2C_{2})^{-1}C_{0}$. Therefore, by (IV.28) and (IV.29) we have
@@
Let $m,n\in\mathbb{N}$ with $2\leq n\leq m$, and let $1\leq q<\infty$. Suppose $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ and $\nabla^{2}f\in L^{n,q}(B^{n})$, then there exist a constant $C>0$ and an open neighborhood $U$ of the origin such that
@@
When $m>n$, Corollary IV.13 is sharp in the sense that there exists $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{m})$ with $\nabla^{2}f\in L^{n,\infty}(B^{n})$ but $f$ is not injective in any neighborhood of the origin. We construct as follows.
Without loss of generality assume $m=n+1$. Define $\tilde{f}_{0}\in C^{\infty}(\mathbb{S}^{n-1},\mathbb{R}^{n})$ by
@@
$f(x)=\begin{dcases}|x|\mkern 2.0mu\tilde{f}\Big(\frac{x}{|x|}\Big),&\text{if }x\neq 0,\\ 0,&\text{otherwise}.\end{dcases}$
Then $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{n+1})$, and there exists a constant $C>0$ such that $|\nabla^{2}f(x)|\leq C|x|^{-1}$ for all $x\in B^{n}\setminus\{0\}$. This implies that $\nabla^{2}f\in L^{n,\infty}(B^{n})$, but $f$ is not injective in any neighborhood of 0.
In the special case $m=n$, a sharp theorem states that any $W^{2,2}$ weak immersion (also called “BLD maps” in this case) is a local homeomorphism everywhere, see [78, Theorem 1.1]. However, when $n\geq 3$, it remains open whether one can replace $W^{2,2}$ by $W^{2,(2,q)}$ for any $2<q<\infty$.
In the proof of [168, Corollary 3.43], it is mentioned that in the case $n=2$, if we assume in addition that $\vec{\Phi}\in W^{2,2}(\Sigma)$, then for any $x\in\mathbb{R}^{m}$, $\theta_{x}$ exists and is in $\mathbb{N}_{0}$. Here we prove a stronger result.
@@
Let $\Sigma$ be an $n$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ with $d\vec{\Phi}\in VMO(\Sigma)$, then we have
$\theta_{y}=\text{Card}(\vec{\Phi}^{-1}(y)),\qquad\text{for all }\,y\in\mathbb{R}^{m}.$ (IV.30)
@@
$\lim_{r\rightarrow 0^{+}}\frac{\int_{f^{-1}(B_{r}(0))}1\,d\textup{vol}_{g_{f}}}{r^{n}|B^{n}|}=1,$ (IV.31)
@@
$f_{r}(w)\coloneqq\frac{f(r\omega)}{r}.$ (IV.32)
@@
$\fint_{B^{n}}\left|\nabla f_{r}-\fint_{B^{n}}\nabla f_{r}\right|=\fint_{B_{r}(0)}\left|\nabla f-\fint_{B_{r}(0)}\nabla f\right|\xrightarrow[r\to 0^{+}]{}0.$ (IV.33)
@@
$\fint_{B^{n}}\nabla f_{C_{0}r_{k_{i}}}\xrightarrow[i\to\infty]{}M.$ (IV.35)
@@
Denote $\tilde{f}_{i}\coloneqq f_{C_{0}r_{k_{i}}}$. By (IV.33), (IV.35) we have
$\lim_{i\rightarrow\infty}\int_{B^{n}}\left|\nabla\tilde{f}_{i}-M\right|=0.$ (IV.36)
Let $\tilde{f}\colon B^{n}\rightarrow\mathbb{R}^{m}$ defined by $\tilde{f}(\omega)=M\omega$. Since $\{\tilde{f}_{i}\}$ is bounded in $W^{1,\infty}(B^{n})$ and $\tilde{f}_{i}(0)=0$ for any $i$, by (IV.36) we obtain that
@@
By (IV.34) and definition of $f_{r}$, for any $\omega\in B^{n}$, $i\in\mathbb{N}$, we have
@@
$|\omega|\leq C_{0}\big|\tilde{f}(\omega)\big|=C_{0}|M\,\omega|.$ (IV.37)
In particular, this implies $M$ has full rank. By (IV.34) we have $f^{-1}\big(B_{r_{k_{i}}}(0)\big)\subset B_{C_{0}r_{k_{i}}}(0)$. We can then apply change of variables formula to get
$(C_{0}r_{k_{i}})^{-n}\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}=\int_{\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)}\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}.$ (IV.38)

### v3

@@ IV.2 Constant of Wente’s inequality for a metric and consequences
Since then, equations that possess this Jacobian structure have been intensively studied. Bethuel–Ghidaglia [20] proved that the constant $C$ is independent of the domain. Ge [63], Braraket [11] and Topping [187] studied such equations on surfaces and proved that the constant is independent of the underlying surface. A weighted Wente inequality has also been studied by Da Lio, Gianocca and Rivière in [46, 65]. In this section, we provide a new proof of the independence of the constant with respect to the underlying metric and domain in Proposition IV.8. As an application, we obtain the existence of a Coulomb frame under the smallness of the Gauss curvature in Proposition IV.10.
@@ IV.2.1 Wente inequality
Let $\Omega\subset\mathbb{R}^{2}$ be a smooth simply connected open set. Let $g$ be a smooth metric on $\Omega$. Let $a,b\in W^{1,2}(\Omega)$ and $u\in W^{1,1}(\Omega)$ be the solution to
$\left\{\begin{aligned} &-\Delta_{g}u=*_{g}(da\wedge db)\quad&&\text{in }\Omega,\\ &u=0\quad&&\text{on }\partial\Omega.\end{aligned}\right$. (IV.16)
@@
  • The first equation of (IV.16) can also be written as
@@
By elliptic regularity, it holds $G_{p}\in C^{\infty}(\Omega\setminus\{p\})$. Thanks to Sard’s theorem (see for instance [111, Chapter 6, Theorem 6.10]), for almost every $\alpha>0$, the set $G_{p}^{-1}(\{\alpha\})$ is a regular curve in $\Omega$. We orient this curve with the vector field
@@
$\int_{\omega_{p}(\alpha,\beta)}|dG_{p}|^{2}_{g}\,d\textup{vol}_{g}=\int_{\alpha}^{\beta}\int_{G_{p}^{-1}(\{s\})}|dG_{p}|_{g}\,d\textup{vol}_{g}\,ds=\beta-\alpha.$ (IV.17)
Now, the solution $u$ to (IV.16) is given by
$u(p)=\int_{\Omega}G_{p}(x)\,da\wedge db.$ (IV.18)
@@
$\leq 2\int_{\omega_{p}(n,n+1)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}.$ (IV.19)
From (IV.18), it holds
@@
$=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$ (IV.20)
$\qquad+\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}.$ (IV.21)
We estimate the term (IV.20) by Cauchy–Schwarz:
@@
$\leq 2\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}.$ (IV.22)
We estimate the term (IV.21) by integration by parts:
@@
$=$ $-\sum_{n\in\mathbb{N}^{*}}\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}.$ (IV.23)
@@
Coming back to (IV.23), we obtain
@@
Using the choice of the $\alpha_{n}$ in (IV.19), we obtain
@@
$\leq 8\,\|da\|_{L^{2}(\Omega,g)}^{2}+8\|db\|_{L^{2}(\Omega,g)}^{2}=16.$ (IV.24)
We used Cauchy–Schwarz and the equalities $\|da\|_{L^{2}(\Omega,g)}=\|db\|_{L^{2}(\Omega,g)}=1$ for the last estimate. Coming back to (IV.18), together with (IV.22) and (IV.24), we deduce that
@@
We multiply (IV.16) by $u$ and integrate by parts:
$\int_{\Omega}|du|^{2}_{g}\,d\textup{vol}_{g}$ $=-\int_{\Omega}u\,\Delta_{g}u\,d\textup{vol}_{g}\leq\|u\|_{L^{\infty}(\Omega)}\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}\leq 18.$
@@ IV.2.2 Estimate of the moving frame
Thanks to [81, Lemma 4.1.3], there exists a frame $(\vec{e}_{1},\vec{e}_{2})\in W^{1,2}(\Omega,g)$ such that
$\begin{cases}d^{*_{g}}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{in }\Omega,\\ \left\langle\partial_{\nu}\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{on }\partial\Omega,\end{cases}$ (IV.25)
where $\nu$ is the unit outward-pointing normal for $g$. Moreover, by decomposing each $d\vec{e}_{i}$ into the sum of its tangential part along $\vec{\Phi}$ and its normal part, we obtain the following estimate
$\int_{\Omega}|d\vec{e}_{1}|^{2}_{g,\mathbb{R}^{m}}+|d\vec{e}_{2}|^{2}_{g,\mathbb{R}^{m}}\,d\textup{vol}_{g}\leq 2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$ (IV.26)
By Poincaré Lemma and (IV.25), there exists $\lambda\in W^{1,2}(\Omega,g)$ such that
$\begin{cases}d\lambda=*_{g}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.27)
Indeed, the boundary condition in (IV.25) implies that $\lambda$ is constant on $\partial\Omega$. Since the first relation of (IV.27) depends only on $d\lambda$ and not on $\lambda$ itself, we can add a constant in order to get the second relation of (IV.27). We apply $d^{*_{g}}$ to (IV.27) to obtain
$\begin{cases}-\Delta_{g}\lambda=-*_{g}d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.28)
@@
Denote $\varepsilon_{0}\coloneqq\|K_{g}\|_{L^{1}(\Omega,g)}$. We multiply the first relation of (IV.28) by $\lambda$ and integrate by parts:
@@ IV.3 Local injectivity and integral densities
In particular, $\nabla f\in VMO(B^{n})$. Then there exists a constant $C>0$ and an open neighborhood $U$ of the origin, both depending only on $\Lambda,\,\omega,\,n,\,m$ such that
@@
$|Mv|\geq C_{0}|v|.$ (IV.29)
Moreover, the constants $C_{0},\,\varepsilon$, and $K$ can be chosen to depend on $\Lambda,\,\omega,\,n,\,m$ only. We choose $r_{0},\varepsilon_{0}\in(0,1/8)$ small enough (to be determined later) depending on $\Lambda,\,\omega,\,n,\,m$ only such that for every ball $B$ of radius $\leq 4r_{0}$ we have $\fint_{B}\left|\nabla f-\fint_{B}\nabla f\right|<\min(\varepsilon,\varepsilon_{0})$. Now fix $x_{0},y_{0}\in B_{r_{0}}(0)$. Let $r\coloneqq|x_{0}-y_{0}|$ and consider the ball $B^{\prime}\coloneqq B_{2r}(x_{0})$. We denote
@@
$\fint_{B^{\prime}}|\nabla f-M_{0}|<\min(\varepsilon,\varepsilon_{0}).$
Hence, it holds $M_{0}\in K^{\prime}$. Letting $f_{0}(x)\coloneqq f(x)-M_{0}x$, we apply Morrey’s inequality [58, Theorem 5.6.4] in the case $p=2n$ and John–Nirenberg inequality [78, Theorem 3.5], for $\varepsilon_{0}$ small enough we have
$\begin{aligned} |f_{0}(x_{0})-f_{0}(y_{0})|&\leq C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f_{0}|^{2n}\right)^{\frac{1}{2n}}\\ &=C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f-M_{0}|^{2n}\right)^{\frac{1}{2n}}\\[5.69054pt] &\leq C_{2}\mkern 2.0mur\mkern 2.0mu\varepsilon_{0}=2^{-1}C_{0}\mkern 2.0mur=2^{-1}C_{0}|x_{0}-y_{0}|,\end{aligned}$ (IV.30)
where $C_{1},\,C_{2}$ are positive constants depending only on $n$, and we choose $\varepsilon_{0}=(2C_{2})^{-1}C_{0}$. Therefore, by (IV.29) and (IV.30) we have
@@
Let $m,n\in\mathbb{N}$ with $2\leq n\leq m$, and let $1\leq q<\infty$. Suppose $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ and $\nabla^{2}f\in L^{n,q}(B^{n})$, then there exists a constant $C>0$ and an open neighborhood $U$ of the origin such that
@@
When $m>n$, Corollary IV.13 is sharp in the sense that there exists $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{m})$ with $\nabla^{2}f\in L^{n,\infty}(B^{n})$ but $f$ is not injective in any neighborhood of the origin. We construct it as follows. Without loss of generality, we can assume $m=n+1$. We define $\tilde{f}_{0}\in C^{\infty}(\mathbb{S}^{n-1},\mathbb{R}^{n})$ by
@@
$f(x)=\begin{dcases}|x|\mkern 2.0mu\tilde{f}\Big(\frac{x}{|x|}\Big)&\text{if }x\neq 0,\\ 0&\text{otherwise}.\end{dcases}$
Then we have $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{n+1})$, and there exists a constant $C>0$ such that $|\nabla^{2}f(x)|\leq C|x|^{-1}$ for all $x\in B^{n}\setminus\{0\}$. This implies that $\nabla^{2}f\in L^{n,\infty}(B^{n})$, but $f$ is not injective in any neighborhood of 0.
In the special case $m=n$, a sharp theorem states that any $W^{2,2}$ weak immersion (also called “BLD maps” in this case) is a local homeomorphism everywhere, see [79, Theorem 1.1]. However, when $n\geq 3$, it remains open whether one can replace $W^{2,2}$ by $W^{2,(2,q)}$ for any $2<q<\infty$.
In the proof of [171, Corollary 3.43], it is mentioned that in the case $n=2$, if we assume in addition that $\vec{\Phi}\in W^{2,2}(\Sigma)$, then for any $x\in\mathbb{R}^{m}$, $\theta_{x}$ exists and is in $\mathbb{N}_{0}$. Here we prove a stronger result.
@@
Let $\Sigma$ be an $n$-dimensional closed smooth manifold and $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ be such that $d\vec{\Phi}\in VMO(\Sigma)$. Then we have
$\theta_{y}=\text{Card}(\vec{\Phi}^{-1}(y)),\qquad\text{for all }\,y\in\mathbb{R}^{m}.$ (IV.31)
@@
$\lim_{r\rightarrow 0^{+}}\frac{\int_{f^{-1}(B_{r}(0))}1\,d\textup{vol}_{g_{f}}}{r^{n}|B^{n}|}=1,$ (IV.32)
@@
$f_{r}(w)\coloneqq\frac{f(r\omega)}{r}.$ (IV.33)
@@
$\fint_{B^{n}}\left|\nabla f_{r}-\fint_{B^{n}}\nabla f_{r}\right|=\fint_{B_{r}(0)}\left|\nabla f-\fint_{B_{r}(0)}\nabla f\right|\xrightarrow[r\to 0^{+}]{}0.$ (IV.34)
@@
$\fint_{B^{n}}\nabla f_{C_{0}r_{k_{i}}}\xrightarrow[i\to\infty]{}M.$ (IV.36)
@@
Denote $\tilde{f}_{i}\coloneqq f_{C_{0}r_{k_{i}}}$. By (IV.34), (IV.36) we have
$\lim_{i\rightarrow\infty}\int_{B^{n}}\left|\nabla\tilde{f}_{i}-M\right|=0.$ (IV.37)
Let $\tilde{f}\colon B^{n}\rightarrow\mathbb{R}^{m}$ defined by $\tilde{f}(\omega)=M\omega$. Since $\{\tilde{f}_{i}\}$ is bounded in $W^{1,\infty}(B^{n})$ and $\tilde{f}_{i}(0)=0$ for any $i$, by (IV.37) we obtain that
@@
By (IV.35) and definition of $f_{r}$, for any $\omega\in B^{n}$, $i\in\mathbb{N}$, we have
@@
$|\omega|\leq C_{0}\big|\tilde{f}(\omega)\big|=C_{0}|M\,\omega|.$ (IV.38)
In particular, this implies $M$ has full rank. By (IV.35) we have $f^{-1}\big(B_{r_{k_{i}}}(0)\big)\subset B_{C_{0}r_{k_{i}}}(0)$. We can then apply the change of variables formula to get
$(C_{0}r_{k_{i}})^{-n}\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}=\int_{\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)}\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}.$ (IV.39)